OpenStudy (anonymous):

Jake is looking over some data regarding the strength, measured in Pascals (Pa), of some building materials and how the strength relates to the length. The data are represented by the exponential function f(x) = 2x, where x is the length. Explain how he can convert this equation to a logarithmic function when strength Pascals.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

x is the length, what is f(x)?

OpenStudy (anonymous):

y?

jimthompson5910 (jim_thompson5910):

in terms of the word problem

OpenStudy (anonymous):

the strength in terms of the length

jimthompson5910 (jim_thompson5910):

yes, x = length, f(x) = strength

jimthompson5910 (jim_thompson5910):

so when they say `strength is 8 Pascals.` we can replace `f(x)` with `8`

OpenStudy (anonymous):

8=2^x

jimthompson5910 (jim_thompson5910):

now convert that to a log equation

OpenStudy (anonymous):

Ok:) is it log 2x =8?

jimthompson5910 (jim_thompson5910):

I think you mean \[\Large \log_2(x) = 8\] right? you're close but not quite there

OpenStudy (anonymous):

yes! Something along those lines

jimthompson5910 (jim_thompson5910):

you'll use the rule \[\Large b^x = y \ \ \implies \ \ \log_b(y) = x\]

OpenStudy (anonymous):

Ok thank-you :)!!

OpenStudy (anonymous):

what about the 8 pascals?

jimthompson5910 (jim_thompson5910):

how would you use that rule to rewrite \[\Large 2^x = 8\]

OpenStudy (anonymous):

Look at base b^x=y as if it were 2^x=8

OpenStudy (anonymous):

Now look at the log equation and replace each variable with the corresponding one.

jimthompson5910 (jim_thompson5910):

yeah so b = 2 and y = 8

OpenStudy (anonymous):

That's it right?

jimthompson5910 (jim_thompson5910):

what log equation do you have now

OpenStudy (anonymous):

2^x = 8 means log(base 2) 8 = x

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

That's it then?

jimthompson5910 (jim_thompson5910):

yeah \[\Large 2^x = 8\] turns into \[\Large \log_2(8) = x\]

OpenStudy (anonymous):

Thank you so much~~!!

OpenStudy (anonymous):

I have one more?

jimthompson5910 (jim_thompson5910):

go ahead

jimthompson5910 (jim_thompson5910):

what are your thoughts?

OpenStudy (anonymous):

it's suppose to be 50^x. Sorry!

OpenStudy (anonymous):

I would tell her to convert it. I dont know which formula to use now though

jimthompson5910 (jim_thompson5910):

you'd use the one I just posted. The rule going from exponential to log

OpenStudy (anonymous):

Ok. Lemme try it

OpenStudy (anonymous):

log 50 (17)=x

jimthompson5910 (jim_thompson5910):

yes, \[\Large \log_{50}(17) = x\]

OpenStudy (anonymous):

That's it? It was that simple?

jimthompson5910 (jim_thompson5910):

here is a pic that has an alternate route http://ashikmdigitalportfolio.weebly.com/uploads/1/9/2/5/19253115/8805020_orig.jpg?146 either method gets the same answer

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

Can I ask one last question? Pleasee?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Shannon manages a small zoo and she has been analyzing the attendance data. Shannon finds that the number of visitors increases exponentially as the temperature increases, and this situation is represented by the function f(x) = 3x. Shannon also finds a linear equation that models the number of people who leave the park early depending on the change in temperature, and it is represented by f(x) = −x + 4. The graph of the two functions is below. Find the solution to the two functions and explain what the solution represents.

OpenStudy (anonymous):

Ty!

jimthompson5910 (jim_thompson5910):

it says `The graph of the two functions is below` where do the two functions cross?

OpenStudy (anonymous):

oh hold on! I can graph it

jimthompson5910 (jim_thompson5910):

looks good

jimthompson5910 (jim_thompson5910):

where do the two functions cross?

OpenStudy (anonymous):

at (1,3)

jimthompson5910 (jim_thompson5910):

what does that point mean?

OpenStudy (anonymous):

It is the solution?

jimthompson5910 (jim_thompson5910):

it's the solution to the system, yes but what does that solution mean in terms of the word problem?

OpenStudy (anonymous):

The number of people who leave the park early due to the temperature?

jimthompson5910 (jim_thompson5910):

let's go back to each function one by one f(x) = 3^x what is x? what is f(x)?

OpenStudy (anonymous):

umm the increase in temperature for f(x)?

jimthompson5910 (jim_thompson5910):

for f(x) = 3^x x = temperature f(x) = number of visitors

jimthompson5910 (jim_thompson5910):

how about g(x) = -x+4 ?

OpenStudy (anonymous):

-x=temperature or people leaving?

jimthompson5910 (jim_thompson5910):

x looks like the change in temp

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hmm..is that when the temperature and # of people are equal?

OpenStudy (anonymous):

Wait! NVM haha

jimthompson5910 (jim_thompson5910):

so what I'm thinking is that when x = 1, the temperature change is 1 degree so as the temp increases by 1, the number of visitors is 3 (maybe 3 thousand or something) also, when x = 1, the number of people who leave early is 3 (thousand?) this problem is a bit odd

OpenStudy (anonymous):

I know right? Ugh!!! So, would (1,3) mean that that's the temp. needed to have the most visitors?

jimthompson5910 (jim_thompson5910):

maybe it's when the temp change is +1, then the number of visitors equals the number of people who leave early

jimthompson5910 (jim_thompson5910):

what's strange is that I don't see how you can get more visitors growing forever if it's like 100+ degrees. If anything, the attendance would go down

OpenStudy (anonymous):

Oh! I didn't look at it that way! That is strange.

OpenStudy (anonymous):

It's a poorly worded question tbh

jimthompson5910 (jim_thompson5910):

yeah I agree

OpenStudy (anonymous):

i'm going to go with your answer :)

OpenStudy (anonymous):

Thanks so much for all your help today!!! You're awesome :D

jimthompson5910 (jim_thompson5910):

sure thing